# SoTA with Less: MCTS-Guided Sample Selection for Data-Efficient Visual Reasoning Self-Improvement

### v2

@@ 4.3 Comparison with Online Data Selection
In this section, we compare our offline data-selection strategy with an online alternative and evaluate their impact on model performance. We adopt an online baseline based on self-consistency filtering: during training we keep only those samples whose rollout accuracy is greater than 0 but below 0.9, drawing additional samples until the training batch is full. Table 4 compares this online variant with our MCTS-based offline selector and a plain offline self-consistency baseline. Similar to the findings in other RL studies [90], the online filter offers negligible improvement except converges faster. The decisive factor is still the ability to identify examples that are truly challenging for the current model, a task at which our MCTS selector excels due to its explicit tree search.
@@ Acknowledgment
Wang and Huang are supported by DARPA Transfer from Imprecise and Abstract Models to Autonomous Technologies (TIAMAT) 80321, National Science Foundation NSF-IIS-2147276 FAI, National Science Foundation NAIRR240045, DOD-AFOSR-Air Force Office of Scientific Research under award number FA9550-23-1-0048.

### v3

@@ 4.3 Comparison with Online Data Selection
In this section, we compare our offline data-selection strategy with an online alternative and evaluate their impact on model performance. We adopt an online baseline based on self-consistency filtering: during training we keep only those samples whose rollout accuracy is greater than 0 but below 0.9, drawing additional samples until the training batch is full. Table 4 compares this online variant with our MCTS-based offline selector and a plain offline self-consistency baseline. Similar to the findings in other RL studies (Yu et al., 2025), the online filter offers negligible improvement except converges faster. The decisive factor is still the ability to identify examples that are truly challenging for the current model, a task at which our MCTS selector excels due to its explicit tree search.
@@ Acknowledgment
Wang and and Huang are supported by DARPA Transfer from Imprecise and Abstract Models to Autonomous Technologies (TIAMAT) 80321, DARPA HR001124S0029-AIQ-FP-019, DOD-AFOSR-Air Force Office of Scientific Research under award number FA9550-23-1-0048, National Science Foundation NSF-IIS-2147276 FAI, National Science Foundation NAIRR240045, National Science Foundation TRAILS Institute (2229885). Private support was provided by Peraton.
Appendix